OpenStudy (asia):

at a kennel, the ratio of dogs to cats was 4 to 9. If there were 64 dogs, how many cats were at the kennel?

OpenStudy (anonymous):

The ratio of dogs to cats is 4 to 9. That means the number of cats divided by the number of dogs is equal to (4/9). So C/D = 4/9 D = 64. Substitute and solve for C.

OpenStudy (radar):

4:9 ::d:c 4:9 :: 64:c 4c=9*64 c =9*16=144 That way you don't have a fraction of a cat